OpenStudy (anonymous):

What is the derivative of 1/x

OpenStudy (anonymous):

(-1/x^2)

OpenStudy (anonymous):

oh thank you!

OpenStudy (anonymous):

\[\frac{1}{x} = x^{-1}\]\[\frac{d}{dx}(x^{-1})= -x^{-2} = \frac{-1}{x^2}\]